Gravitational Fields: a) What is the escape velocity for a stationary object (of mass 100 kg) from a planet if an object dropped on its surface falls 5m in 1.4 seconds. (diameter of planet is 12,000 km) b) What is the objects velocity at a distance of 60, 000 km if the velocity is 1.2 times the escape velocity?

Please show and explain as much work as possible. Thank you!
Answer #1

Ans: a) Calculating g. I got s = ut ti (2nd Eqr of an of Motor S o + 1 xgxce) ... (as it is dropped IS.102 m/s2 g= veloc cityvesca 12x5-102X 6000x103 esca =78221.77m/s Ans.b v1.2 vesc = 1-2 X 7824.577 2 9389.4924m/s)